3 hours to drive 135 miles you drive 55 miles per hour on highways and 40 miles per hour the remaining time how much time did you spend driving each speed?
Answer by palanisamy(496)

Let the time taken to drive 55 miles per hour on highways =x
Let the time taken to drive 40 miles per hour =y
Then, total time x+y=3 ...(1)
Total distance 55x+40y=135 ...(2)
(1)*40=> 40x+40y=120 ...(3)
(2)-(3)=> 15x=15
x=1
y=3-1=2